# A Bayesian Interpretation of Adaptive Low-Rank Adaptation

### v1

@@ I Introduction
Motivated by AdaLoRA, we are primarily interested in the importance scoring mechanism as it can be generically applied to PEFT for parameter selection. One of the main criticisms of AdaLoRA is that it uses a sensitivity-based importance score which lacks a rigorous theoretical foundation. [4] However, there are importance metrics with strong theoretical support, many of which originate from Bayesian neural networks (BNNs). A widely recognized metric is the signal-to-noise ratio (SNR) [5, 6, 7], commonly used in BNN pruning and compression. The interpretation is straightforward: a low SNR makes the neuron’s output too noisy to be useful, while a high SNR indicates valuable, low-noise output. The SNR could be a drop-in replacement for the sensitivity for AdaLoRA, allowing the pruning of SVD triplets with low SNRs during fine-tuning for dynamic rank adjustment.
@@
In this study, we leverage Bayesian importance metrics alongside the Improved Variational Online Newton (IVON) optimizer, a recent advance in VI, to develop a Bayesian counterpart to AdaLoRA, utilizing SNR as the importance score. By comparing its performance with the sensitivity-based importance metric on the GLUE benchmark [10], we demonstrate that the Bayesian approach not only achieves comparable or superior performance but also offers a 10% speed-up over the original AdaLoRA with Adam. A closer examination of the underlying theory reveals a strong connection between these two metrics, providing a Bayesian interpretation of the sensitivity as an importance score. Additionally, our findings indicate that the magnitude, rather than the variance, is the primary indicator of the importance of parameters. The source code is provided. [^1] https://github.com/idiap/vilora
@@ II-C Bayesian Importance Scores
Despite the empirical success of the sensitivity-based importance metric, there still lacks an underlying theory explaining its effectiveness. Given the criticism, we focus on theoretically supported importance metrics, many of which originate from Bayesian neural networks (BNN). BNNs model weights as probability distributions, enabling the network to quantify uncertainties in its predictions. The most commonly used distribution is the Gaussian distribution, therefore the model is parameterized by two sets of parameters: the mean $\mu$ and the standard deviation $\sigma$ (or the variance $\sigma^{2}$).
@@ III Variational Inference
The Improved Variational Online Newton (IVON) [8] is a recent VI optimizer that matches the performance of Adam at a comparable computational cost. Its key innovations include bypassing the expensive per-example gradient square computation through a reparameterization trick and incorporating several practical techniques to enhance performance. IVON stands out as the first VI optimizer proven to be both effective and efficient for training large networks, while still delivering the benefits of VI. In our experiments, we utilize IVON to estimate parameter variance, enabling the use of SNR as an importance metric.
@@ IV-B Implementation Details
We base our code on the text classification examples of the Hugging Face Transformers library [20] and the Parameter-Efficient Fine-Tuning (PEFT) library [21]. For IVON, we use the official implementation [^2] https://github.com/team-approx-bayes/ivon. We compare the methods under two budget configurations where the target rank is set to 2 and 4 respectively, resulting in the total trainable parameters being 0.3M and 0.6M (of 183M). Full fine-tuning and LoRA applied to all modules are also added as baselines.
@@ IV-C Training and Evaluation
Figure 1: Comparison of rank distributions after fine-tuning on MNLI, with deeper colors indicating higher ranks. Results are averaged across five runs with different random seeds. $W_{q}$, $W_{k}$, $W_{v}$, $W_{o}$: weights of the query, key, value, output layers of attention; $W_{f_{1}}$, $W_{f_{2}}$: weights of the feed-forward layers.
@@ V Results and Analyses
The main results are shown in Table I. Note that we sort the tasks according to dataset sizes and divide them into two groups since we notice that IVON needs extra tricks to ensure good results on small datasets. In general, all PEFT methods outperform full fine-tuning, and AdaLoRA outperforms LoRA. Switching the optimizer from Adam to IVON results in comparable performance, demonstrating that IVON is capable of state-of-the-art performance in PEFT. We further elaborate our findings from the following perspectives.
@@ V-D A Bayesian Interpretation of Sensitivity
The similarity in performance and the rank distribution between the sensitivity and the SNR suggests a close relationship between them. A closer examination of the underlying theory reveals that sensitivity is, in fact, aligned with the principles of SNR. Specifically, in IVON, the standard deviation $\sigma$ is estimated as $\sigma=1/\sqrt{\lambda(\mathbf{h}+\delta)}$, where $\mathbf{h}$ is the diagonal Hessian, $\lambda$ is the effective sample size, and $\delta$ is a weight decay term. Notably, $\mathbf{h}$ can be approximated by the expected square gradient on the training data [22], $\mathbf{h}\approx\mathbb{E}_{\mathcal{D}}[(\nabla_{\boldsymbol{\theta}}\ell)^{2}]$. Consequently, the inverse of the standard deviation, $1/\sigma$, in the context of SNR, is proportional to the root mean square of the gradient $\sqrt{\mathbb{E}_{\mathcal{D}}[(\nabla_{\boldsymbol{\theta}}\ell)^{2}]}$, and therefore analogous to the magnitude of the gradient $|\nabla_{\boldsymbol{\theta}}\ell|$. This implies that the sensitivity $|\theta\nabla_{\theta}\ell|$ has the component $|\nabla_{\theta}\ell|$ acting as an uncertainty measure analogous to $1/\sigma$ in SNR, thereby providing a Bayesian interpretation of the sensitivity as an importance metric.
@@ VI Conclusions
In this study, we developed a Bayesian alternative to AdaLoRA, leveraging the signal-to-noise ratio as the importance score with the IVON optimizer. By comparing the performance of different importance metrics, we demonstrated that this Bayesian approach not only matched or surpassed the performance of using the sensitivity-based importance metric on the GLUE benchmark, but also achieved a 10% speed-up over the original AdaLoRA with Adam. The theoretical analysis uncovered a significant link between these two metrics, offering a Bayesian perspective on the efficacy of the heuristic sensitivity-based metric as an importance score. Furthermore, our results suggested that the magnitude, rather than the variance, served as the key indicator of the importance of parameters.
Acknowledgment
This work received funding under NAST: Neural Architectures for Speech Technology, Swiss National Science Foundation grant 185010.

### v2

@@ I Introduction
Motivated by AdaLoRA, we are primarily interested in the importance scoring mechanism as it can be generically applied to PEFT for parameter selection. The sensitivity-based importance metric is originally based on the heuristic that the importance of parameters can be quantified by the error induced by removing them, which in turn can be approximated by the square of the gradient-weight product [4, 5]. Meanwhile, there are importance metrics with strong theoretical support, many of which originate from Bayesian neural networks (BNNs). A widely recognized metric is the signal-to-noise ratio (SNR) [6, 7, 8], commonly used in BNN pruning and compression. The interpretation is straightforward: a low SNR makes the neuron’s output too noisy to be useful, while a high SNR indicates valuable, low-noise output. The SNR could be a drop-in replacement for the sensitivity-based importance score in AdaLoRA, allowing the pruning of SVD triplets with low SNRs during fine-tuning for dynamic rank adjustment.
@@
In this study, we leverage Bayesian importance metrics alongside the Improved Variational Online Newton (IVON) optimizer, a recent advance in VI, to develop a Bayesian counterpart to AdaLoRA, utilizing SNR as the importance score. By comparing its performance with the sensitivity-based importance metric on the GLUE benchmark [11], we demonstrate that the Bayesian approach not only achieves comparable or superior performance but also offers a 10% speed-up over the original AdaLoRA with Adam. A closer examination of the underlying theory reveals a strong connection between these two metrics, providing a Bayesian interpretation of the sensitivity as an importance score. Additionally, our findings indicate that the magnitude, rather than the variance, is the primary indicator of the importance of parameters. The source code is available. [^1] https://github.com/idiap/vilora
@@ II-C Bayesian Importance Scores
In this work, we focus on theoretically supported importance metrics that originate from Bayesian neural networks (BNN). BNNs model weights as probability distributions, enabling the network to quantify uncertainties in its predictions. The most commonly used distribution is the Gaussian distribution, therefore the model is parameterized by two sets of parameters: the mean $\mu$ and the standard deviation $\sigma$ (or the variance $\sigma^{2}$, we also refer to $\sigma$ as variance for the sake of simplicity).
@@ III Variational Inference
The Improved Variational Online Newton (IVON) [9] is a recent VI optimizer that matches the performance of Adam at a comparable computational cost. Its key innovations include bypassing the expensive per-example gradient square computation through a reparameterization trick and incorporating several practical techniques to enhance performance. IVON stands out as the first VI optimizer proven to be both effective and efficient for training large networks, while still delivering the benefits of VI. In our experiments, we utilize IVON to estimate the variance of parameters, enabling the use of SNR as an importance metric.
@@ IV-B Implementation Details
We base our code on the text classification examples of the Hugging Face Transformers library [22] and the Parameter-Efficient Fine-Tuning (PEFT) library [23]. For IVON, we use the official implementation [^2] https://github.com/team-approx-bayes/ivon. We compare the methods under two budget configurations where the target rank is set to 2 and 4 respectively, resulting in the total trainable parameters being 0.3M and 0.6M (of 86M). Full fine-tuning and LoRA applied to all modules are also added as baselines.
@@ IV-C Training and Evaluation
Figure 1: Comparison of rank distributions after fine-tuning DeBERTaV3-base on MNLI, with deeper colors indicating higher ranks. Results are averaged across five runs with different random seeds. $W_{q}$, $W_{k}$, $W_{v}$, $W_{o}$: weights of the query, key, value, output layers of attention; $W_{f_{1}}$, $W_{f_{2}}$: weights of the feed-forward layers.
@@ V Results and Analyses
The main results are shown in Table I. For MNLI, the “matched” validation set was used for evaluation. Note that we sort the tasks according to dataset sizes and divide them into two groups since we notice that IVON needs extra tricks to ensure good results on small datasets. In general, all PEFT methods outperform full fine-tuning, and AdaLoRA outperforms LoRA. Switching the optimizer from Adam to IVON results in comparable performance, demonstrating that IVON is capable of state-of-the-art performance in PEFT. We further elaborate our findings from the following perspectives.
@@ V-D A Bayesian Interpretation of Sensitivity
The similarity in performance and the rank distribution between the sensitivity and the SNR suggests a close relationship between them. A closer examination of the underlying theory reveals that sensitivity is, in fact, aligned with the principles of SNR. Specifically, in IVON, the standard deviation $\boldsymbol{\sigma}$ is calculated as $\boldsymbol{\sigma}=1/\sqrt{\lambda(\mathbf{h}+\delta)}$, where $\mathbf{h}$ is the diagonal Hessian, $\lambda$ is the effective sample size, and $\delta$ is a weight decay term. Notably, $\mathbf{h}$ can be approximated by the expected square gradient on the training data [24], $\mathbf{h}\approx\mathbb{E}_{\mathcal{D}}[(\nabla_{\boldsymbol{\theta}}\ell)^{2}]$, also known as the diagonal of the expected Fisher information matrix (FIM). Consequently, the inverse of the standard deviation, $1/\boldsymbol{\sigma}$, in the context of SNR, is akin to the root mean square of the gradient $\sqrt{\mathbb{E}_{\mathcal{D}}[(\nabla_{\boldsymbol{\theta}}\ell)^{2}]}$, and therefore analogous to the magnitude of the gradient $|\nabla_{\boldsymbol{\theta}}\ell|$. This implies that the sensitivity $|\theta\nabla_{\theta}\ell|$ has the component $|\nabla_{\theta}\ell|$ acting as an uncertainty measure analogous to $1/\sigma$ in SNR, thereby providing a Bayesian interpretation of the sensitivity as an importance metric. These findings resonate with the comment in [5] that the sensitivity has connections with the FIM. Note that both methods adopt exponential moving average smoothing to compute the global value of the corresponding metric during training. The main difference is that the smoothing is applied to the magnitude of the gradient-weight product in AdaLoRA, while the SNR is computed using the global Hessian tracked by IVON.
@@ VI Conclusions
In this study, we developed a Bayesian alternative to AdaLoRA, leveraging the signal-to-noise ratio as the importance score with the IVON optimizer. By comparing the performance of different importance metrics, we demonstrated that this Bayesian approach not only matched or surpassed the performance of using the sensitivity-based importance metric on the GLUE benchmark, but was also a faster alternative to the original AdaLoRA with Adam. The theoretical analysis uncovered a significant link between these two metrics, offering a Bayesian perspective on the efficacy of the heuristic sensitivity-based metric as an importance score. Furthermore, our results suggested that the magnitude, rather than the variance, served as the key indicator of the importance of parameters.